What is meant by transfer function of a system


1
Expert's answer
2021-01-21T09:56:14-0500

Answer

  • A transfer function is an intermediate function that comprises & explains the theoretical procedures on how the input is transferred to the relevant output.
  • It is applicable from simple algebraic sums, engineering applications to everyday phenomena.


Examples

  • In "\small Y=\large\frac{a-b}{2}\times x", the input "\small x" is transfered to its output "\small Y" according to the transfer function "\small \frac{a-b}{2}"
  • How water at room temperature boils upon the received heat is the transfer function of that phenomenon.


  • But a transfer function is usually described only for engineering applications such as control systems & yet it is involved in every cause & result system.